A perfect gas at 27 degree celcius is heated at constant pressure so as to triple its volume The temp of gas will be

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear Pallavi

T1 = 27 +273 =300 K

V1 =V

V2=3V

T2 =?

use  P1V1/T1  = P2V2/T2

GIVEN P1 =P2

 SO V1/T1  =V2/T2

V/300 = 3V/T2

 T2 =900K

= 627 C

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

All the best.
Regards,
Askiitians Experts
Badiuddin

Indar Preet Kainth
6 years ago

Dear Pallavi

T1 = 27 +273 =300 K 

V1 =V

V2=2V 

T2 =?

use  P1V1/T1  = P2V2/T2 

GIVEN P1 =P2

 SO V1/T1  =V2/T2 

V/300 = 2V/T2

 T2 =600K 

= 327 C

We are not IITians, we are ground technicians.


All the best.
Regards,
Inderpreet Singh

Kushagra Madhukar
2 years ago
Dear student,
Please find the solution to your problem.
 
T1 = 27 +273 =300 K
V1 =V
V2=3V
T2 =?
use  P1V1/T1  = P2V2/T2
GIVEN P1 =P2
SO V1/T1  =V2/T2
V/300 = 3V/T2
T2 = 900K = 627oC
 
Thanks and regards,
Kushagra